# CAP[Uncaptioned image] : Evaluation of Persuasive and Creative Image Generation

### v1

@@ 5.2 Creativity
Image Commercial Ads PSA Ads
$I_{input-text}$ T2I $C_{obj}$ $C_{LLM}$ PA $P_{LLM}$ $C_{obj}$ $C_{LLM}$ PA $P_{LLM}$
$I_{AR}$ SDXL 2.03 0.63 0.62 0.60 1.33 0.59 0.48 0.52
AuraFlow 2.12 0.73 0.64 0.60 1.36 0.72 0.42 0.57
$I_{LLM}$ AuraFlow 2.25 0.77 0.70 0.76 1.87 0.71 0.60 0.67
$I_{real}$ 2.28 0.67 0.98 0.66 2.04 0.65 0.60 0.61
Table 4: Comparison of persuasiveness and creativity proposed ($C_{obj}$, PA) and baseline metrics. Best gen. image result bolded.
Annotators Commercial PSA All
H, $C_{LLM}$0.04 0.15 0.04
H, $C_{obj}$ (ours) 0.58 0.53 0.55
@@
Annotators Commercial PSA All
H, $P_{LLM}$ 0.27 0.25 0.26
H, $PC$ 0.83 0.54 0.65
H, $PA$ (ours) 0.94 0.75 0.85
H1, H2 0.89 0.50 0.71
Table 6: Agreement among human annotators, PA score, baseline. PC is the average of scores for PA components without alignment score.
Reliability of the proposed Creativity score. $I_{real}$ are images from PittAd dataset [18] designed to be creative and persuasive and are expected to be more creative than $I_{gen}$. As observed in Table 4 the $C_{LLM}$ score for $I_{real}$, which represents the ground-truth images, is lower than the scores for both $I_{LLM}$ and $I_{AR}$ generated by AuraFlow-2, by 10% and 6% respectively, across both Commercial and PSA advertisements. This raises questions about the reliability of LLMs in accurately scoring creativity, as real images receive lower scores than generated ones. In contrast, our proposed metrics assign the highest scores to $I_{real}$, demonstrating a more consistent alignment with human expectations. As observed in Table 6, the Krippendorff’s Alpha agreement between human annotators and the baseline metric $C_{LLM}$ is -0.04 for Commercial advertisements, indicating systematic disagreement [28], and 0.15 for PSA advertisements, suggesting a low and close to random level of agreement [28]. Our proposed metric, $C_{obj}$, shows significant improvement over these baseline measures, achieving agreement of 0.58 for Commercial advertisements and 0.53 for PSA advertisements, representing moderate agreement [28]. While $C_{obj}$ does not yet reach human-human agreement levels, it significantly decreases the gap, offering a reliable and robust metric for creativity evaluation.
T2I models lack creativity. Table 1, and 4 represents that creativity score drops from $I_{real}$ to $I_{AR}$ by 0.25 and 0.71 for commercial and PSA advertisements, respectively, showing the lack of creativity in T2I models. The bigger gap in PSA advertisements shows that when there are visual objects in the text prompt it is easier for the model to generate a creative scene than when the text input is completely implicit. Our proposed approach for generating creative images is to first generate creative descriptions for the image using the creativity of LLMs and generating an image for the descriptions. Table 1 demonstrates that $I_{LLM}$ achieves higher creativity scores than $I_{AR}$, with increases of 0.22 and 0.54 for commercial and PSA advertisements, respectively, showcasing the effectiveness of using $D_{LLM}$ descriptions for more creative outputs.
@@ 5.3 Persuasiveness
Accuracy of PA. Due to the lack of baseline metrics for persuasiveness, we compare our proposed metric with $P_{LLM}$. As shown in Table 4, $P_{LLM}$ scores $I_{real}$ (ground-truth images) lower than $I_{LLM}$ by 10% and 6% for Commercial and PSA advertisements, respectively, suggesting that LLMs (given image description) may lack the capability to accurately understand and evaluate the persuasiveness of images when asked a single direct question about persuasiveness. In contrast, our proposed metric consistently scores all generated images lower than or equal to the ground-truth images, $I_{real}$, aligning with human expectations. Table 6 represents the agreement levels, where $P_{LLM}$ achieves only minimal agreement with human annotations, at 0.27 (fair agreement [28]) and 0.25 for Commercial and PSA ads, respectively. In contrast, our proposed metric shows a high agreement with human ratings of 0.94 (almost perfect [28]) and 0.75 (substantial agreement [28]), surpassing even the agreement between human annotators.
@@
Table 7: Agreement among annotators and the LLM responses for each persuasiveness component (column headers) used in PA (denoted PA*). “All” computes agreement across all components: for H, we get the image chosen most frequently across all the components, and for PA, we choose the image with higher average score.
@@
Table 8: Scoring different components of persuasiveness.
Accuracy of different components of persuasiveness. Table 7 represents the agreement among human annotators and scores from PA across each component used in the PA score (Fig. 3). When comparing the agreement among human annotators across all images, Tab. 6 and 7 show higher agreement when we directly ask the annotator to choose the more persuasive image (0.71, Tab. 6) compared to the average agreement across individual components (0.52, average of Tab. 7 columns except “All”). This suggests that the individual component questions are more subjective than directly asking the human annotators to choose the more persuasive images. However, average human agreement over these persuasiveness components (Tab. 7) is still reasonable (0.52 is interpreted as moderate agreement), showing the potential for using these questions. Importantly, using these components in formulating the PA prediction of the AI system (through the PA score) is also more helpful than directly asking the LLM to score persuasiveness ($P_{LLM}$). In particular, in Tab. 6, agreement between PC, a variant of PA which averages the PA components without the alignment score CITE, still achieves higher agreement with humans than the single-question $P_{LLM}$.
Low persuasion in images generated by T2I models. As observed in Table 1 , the persuasiveness (PA score) drops by 0.34 and 0.18 from $I_{real}$ to $I_{AR}$ for commercial and PSA advertisements, respectively, indicating a lack of persuasion in images generated by T2I models. Table 1 also shows an increase of 0.06 and 0.18 in PA score from $I_{AR}$ to our proposed approach ($I_{LLM}$). Table 8 compares different images across each component of persuasiveness. $I_{LLM}$ generated by Auraflow-2 outperforms $I_{AR}$ also generated by Auraflow-2 in 6 out of 7 components in both Commercial and PSA advertisements. This further highlights the effectiveness of our proposed approach in generating more persuasive images.
5.4 Comparison of Commercial and PSA ads
Comparing agreement of different alignment metrics with human annotation in Tab. 3 shows that baseline metrics struggle more in the evaluation of PSA advertisements because the prompt is much more implicit than the prompt in Commercial ads and lack detailed visual objects. Comparison of CITE scores for different images in Tab. 2 across PSA and Commercial Advertisement demonstrates that the performance of T2I models drops with decreasing the explicitness of visual scenes and objects in the text input.
Unlike alignment, in creativity, Tab. 6 represents lower agreement between $C_{LLM}$ and human annotators for Commercial advertisements compared to PSA. Due to the presence of visual elements in Commercial advertisement messages, the generated images typically include these elements. However, simply incorporating visual elements does not necessarily make an image appear creative to human annotators. In contrast, the presence of these elements in the description can make the images more complex than text-only images which seems more creative from the perspective of the LLM. In Table 1, $C_{obj}$ drops by 0.71 from $I_{real}$ to $I_{AR}$ for PSAs, while this gap is only 0.25 for Commercial advertisements. The absence of distinct visual elements in PSAs often leads to a higher number of text-only images, contributing to a lower creativity for PSA ads. Table 1 represents higher increase in $C_{obj}$ from $I_{AR}$ to $I_{LLM}$ for PSAs (0.54) compared to Commercial ads (0.22), highlighting the effectiveness of our proposed approach in handling higher levels of implicitness in PSA ads.
@@ 6 Conclusion
We addressed the task of Advertisement Image Generation, focusing on the challenges of generating and evaluating persuasive, and creative advertisement images from visually implicit messages. We introduced three metrics—Creativity, Alignment, and Persuasiveness (CAP)—that achieved high Krippendorff’s Alpha agreement with human annotations, demonstrating their reliability. Our experiments reveal that existing T2I models struggle to generate images that are creative, aligned, and persuasive. To improve performance in this task, we used a simple yet effective approach using LLMs to generate creative and persuasive descriptive prompts from abstract messages. We verified our contributions through agreement with human annotations and CAP metric evaluations, showing the potential of our method to enhance T2I models for advertisement image generation, particularly in cases requiring interpretation of implicit content.
@@
    A detailed table for text-image alignment, including CLIP-scores (Table 9).
@@
  3. 3.
@@ 7 Alignment Results
Image Commercial Ads PSA Ads
$I_{input-text}$ T2I CITE IR VQA CS CITE IR VQA CS
@@
Table 9: Comparison of alignment of the images with action-reason message ($AR_{m}$) using our proposed CITE and baseline metrics. Best result per column bolded (including real images). IR = Image Reward. CS = CLIP-score.
Table 9, represents the text-image alignment scores for different images and evaluated by CITE and baseline metrics. We included most results from this table in the main text, but had omitted Clip-score due to a time constraint. The table highlights that Image-Reward [50] and VQA [25] assign lower scores to $I_{real}$ (ground-truth) compared to generated images. Clip-score [14] performs reasonably as it rates $I_{real}$ the highest, but it rates the alignment of $I_{LLM}$ and $I_{AR}$ equally. This is problematic because human annotators preferred $I_{LLM}$ in 92% of comparisons, demonstrating its superior alignment. CITE assigns the highest score to $I_{real}$, followed by $I_{LLM}$ which highlights the higher accuracy of our proposed evaluation method. Additionally, the agreement between human annotators and each of the alignment metrics, as shown in Table 3 of the main paper, demonstrates the superior accuracy of our proposed metric. Specifically, it outperforms Image-Reward, VQA, and CLIP-score by margins of 0.6, 0.4, and 0.54 out of 1, respectively, highlighting its effectiveness in evaluating alignment.
@@ 8 Persuasiveness Metrics on Real Ads
Due to the inherent differences in quality between $I_{real}$ and $I_{gen}$, we avoided direct comparisons in human evaluations to reduce bias in selecting the better option. Instead, we compared two $I_{real}$ images from the same topic. Since the images were distinct, their corresponding $AR_{m}$ messages also differed, making alignment comparisons non-applicable. However, we evaluated the persuasiveness components for the two $I_{real}$ images and reported the agreement levels between the annotators, as well as between the annotators and the LLM, across all components. The results are presented in Table 10. We also analyze the agreement between the most frequently chosen images by annotators across all components and the average scores assigned to each image for these components. As observed in Table 10, while the agreement for individual components is lower than chance, combining all components results in an average agreement increase of 0.44. It is worth noting that the agreement between the PA* metric and human annotators is higher than the agreement between the human annotators.
@@
Table 10: Agreement among annotators and the LLM responses for each persuasiveness component (column headers) used in PA (denoted PA*). “All” computes agreement across all components: for H, we get the image chosen most frequently across all the components, and for PA, we choose the image with higher average score.
@@ 9 Failure Examples and Limitations
Figure 5: Example of failure of CITE in both generating accurate $AR_{gen}$ and choosing the more aligned image with $AR_{m}$. Red text indicates inaccurate action/reason considering the image. Green text shows accurate action/reason. Red circle shows inaccurate choice of image for which image is more aligned with the $AR_{m}$ according to the human annotation.
In this section, we first highlight examples of failures in the CITE method, specifically in generating accurate $AR_{gen}$ and selecting images that align with human annotators’ choices, as illustrated in Fig. 5. Next we discuss the limitation of our proposed metrics.
@@ 10 Prompts
In this section, we present the prompts used in different evaluation methods. We begin with the prompts for generating descriptions, as shown in Listings LABEL:prompt:train_description and LABEL:prompt:inference_description. Next, we provide the prompts for generating action-reason statements based on image descriptions in Listings LABEL:prompt:alignment_zeroshot and LABEL:prompt:alignment_ft. Then, we present the prompts used for evaluating the creativity and persuasiveness of images with LLMs, detailed in Listings LABEL:prompt:creativity and LABEL:prompt:persuasiveness. Finally, we outline the prompts for evaluating individual PA components, including elaboration, synthesis, originality, imagination, audience targeting, benefit conversion, and appeal category, in Listings LABEL:prompt:elaboration, LABEL:prompt:synthesis, LABEL:prompt:originality, LABEL:prompt:imagination, LABEL:prompt:audience, LABEL:prompt:benefit, and LABEL:prompt:appeal.
@@
Listing 2: Description generation for images using InternVL-v2-26B[7] prompt templates for inference. If the answer to Q1 is “No”, then the CITE score for that image is 0.
@@
Listing 3: Prompt for Action-Reason Prediction with zero-shot LLAMA3-Instruct[9]. {Blue} denotes elements added dynamically, based on the description of the image.
@@
Listing 4: Prompt for Action-Reason Prediction with fine-tuned LLAMA3-Instruct[9] in CITE evaluation. {Blue} denotes elements added dynamically, based on the description of the image.
@@
Listing 5: Prompt for evaluation of creativity with LLAMA3-Instruct[9]. {Blue} denotes elements added dynamically, based on the description of the image.
@@
Context: If the image convinces the audience to take an action it is considered persuasive.
@@
Question: Based on the context score the persuasiveness of the described image in range of (0, 5).
@@
Listing 7: Prompt for scoring elaboration with LLAMA3-Instruct[9]. {Blue} denotes elements added dynamically.
@@
How visually detailed the image is? Do not consider the text in the image. Return a score in range (0,5).
@@
Listing 8: Prompt for scoring synthesis with LLAMA3-Instruct[9]. {Blue} denotes elements added dynamically.
@@
How well the image connects the objects that are usually unrelated? Return the score in range (0,5).
@@
Listing 9: Prompt for scoring originality with LLAMA3-Instruct[9]. {Blue} denotes elements added dynamically.
@@
Listing 10: Prompt for scoring imagination with LLAMA3-Instruct[9]. {Blue} denotes elements added dynamically.
@@
Assume you are a human evaluating the given image. How well the image allowed you to form images you have not directly experienced before more easily? Do not consider the text in the image. Return a score in range (0,5)
@@
Listing 11: Prompt for scoring how well the image targets audience with LLAMA3-Instruct[9]. {Blue} denotes elements added dynamically.
@@
Listing 12: Prompt for scoring how well the image converts features to benefits for costumer with LLAMA3-Instruct[9]. {Blue} denotes elements added dynamically.
@@
How well the image connects the feature of products to the benefits for costumers? Return the score in range (0,5).
@@
Listing 13: Prompt for scoring how well the image targets correct appeal category with LLAMA3-Instruct[9]. {Blue} denotes elements added dynamically.
@@
The are three types of rhetorical appeals or ways to convince the audience:
@@
Question: Assume you are a human evaluating the image. Based on the context score how well the image appeals to {appealcategory} in range of (0,5).

### v2

@@ 5.2 Creativity
Figure 5: Creativity: (d) which shows unique but relevant object portrayal scores best, while random (a) and generic objects (b) score worst. Red/Green shows low/high values for $C_{obj}$, low/high for $AIM$, & high/low for SimSim. Blue denotes moderate values.
Qualitative example. Fig. 5 shows examples of $C_{obj}$ and its components AIM and Sim. Fig. 5a was generated based on $AR_{m}$ “I should shop here because they have good sales” (objects: sales, foods, deals). It shows unusual objects (good Sim score) which are however irrelevant (bad AIM score), hence bad creativity overall. Fig. 5b (“I should drink Coca-Cola because attractive women do”, obj: Coca-Cola) achieves the lowest score (moderate AIM, typical objects). Fig. 5c has high alignment with the prompt, but the objects are not creative. Fig. 5d (“I should drink this root bear because it has a discount”; obj: root bear, A&\&W drink) has the highest overall score due to unusual portrayal of the objects and good alignment with the prompt.
Annotators COM PSA All
H, $C_{LLM}$ -0.03 0.15 0.04
H, $C_{obj}$ 0.57 0.53 0.54
@@
Reliability of the proposed Creativity score. As observed in Table 3, the Krippendorff’s Alpha agreement between human annotators and the baseline metric $C_{LLM}$ is -0.03 for Commercial advertisements, indicating systematic disagreement [31], and 0.15 for PSA advertisements, suggesting a low and close to random level of agreement [31]. Our proposed metric, $C_{obj}$, shows significant improvement over these baseline measures, achieving agreement of 0.57 for Commercial and 0.53 for PSA advertisements, representing moderate agreement [31]. While $C_{obj}$ does not yet reach human-human agreement levels, it significantly decreases the gap, offering a reliable and robust metric for creativity evaluation. In supp, we include results from baseline metrics on both real and generated ads.
T2I models lack creativity. In Table 2, we observe $I_{LLM}$ achieves higher creativity scores than $I_{AR}$ in all cases, showcasing the effectiveness of using $\mathcal{E}_{LLM}$ expansions. QwenLM achieves significantly better results than LLAMA3 as the source of the expansion. Among LLMs for evaluation (AIM), LLAMA3 has a slight edge for commercial ads, while other scores are similar.
Annotators COM PSA All
H, $P_{LLM}$ 0.27 0.26 0.27
H, $P_{comp}$ (InternVL, LLAMA3) 0.83 0.54 0.65
H, $P_{comp+AIM}$ (InternVL, LLAMA3) 0.85 0.75 0.80
H, $P_{comp+AIM}$ (QwenVL, LLAMA3) 0.73 0.63 0.68
H, $P_{comp+AIM}$ (InternVL, QwenLM) 0.89 0.30 0.63
H, $P_{comp+AIM}$ (QwenVL, QwenLM) 0.89 0.74 0.74
H1, H2 0.80 0.56 0.70
Table 4: Agreement with human annotators for different persuasiveness scores. We show different combinations of MLLMs and LLMs for $P_{comp+AIM}$ by ([MLLM], [LLM]).
@@ 5.3 Persuasiveness
Accuracy of $P_{comp+AIM}$. Due to the lack of baseline metrics for persuasiveness, we compare our proposed metric with $P_{LLM}$, in Table 4. All $P_{comp}$ and $P_{comp+AIM}$ results outperform $P_{LLM}$, confirming the value of our components-based persuasion evaluation. Further, the difference between $P_{comp+AIM}$ and $P_{comp}$ shows the value of adding alignment to the components.
@@
Table 5: Agreement among annotators and LLM responses for each persuasiveness component (column headers). “All” compares the image chosen most frequently across the components by humans, and the image with higher average score from $P_{comp+AIM}$.
@@
Table 6: Scoring direct and expansion-based T2I generation (and real ads) using different components of persuasiveness.
Accuracy of different components of persuasiveness. Table 5 represents the agreement among human annotators and scores from $P_{comp}$ across each component used in the $P_{comp}$ score (Fig. 3). When comparing the agreement among human annotators across all images, Tab. 4 and 5 show higher agreement when we directly ask the annotator to choose the more persuasive image (0.80, Tab. 4) compared to agreement with combined per-component responses (0.78 under “All”, Tab. 5). This suggests that the individual component questions are more subjective than directly asking the human annotators to choose the more persuasive images. However, average human agreement over these persuasiveness components (Tab. 5) is still reasonable: 0.53 average of Tab. 5 columns except “All”, is interpreted as moderate agreement, and the combined questions (“All”) agreement is 0.89, showing the potential for using these questions. Importantly, using these components in formulating the $P_{comp+AIM}$ prediction of the AI system is also more helpful than directly asking the LLM to score persuasiveness ($P_{LLM}$, Tab. 4).
Low persuasion in images generated by T2I models. As observed in Table 2, the best $P_{comp+AIM}$ scores are always achieved using LLM expansion. Evaluation using QwenVL as the MLLM is slightly worse than InternVL. Among T2I methods, FLUX is worse than AuraFlow (top part of the table). Table 6 compares different images across each component of persuasiveness. $I_{LLM}$ (using LLAMA3) generated by Auraflow outperforms $I_{AR}$ also generated by Auraflow in 6 out of 7 components in both Commercial and PSA advertisements. This further highlights the effectiveness of our proposed approach in generating more persuasive images.
@@ 6 Conclusion
We examined the challenges of generating and evaluating persuasive and creative advertisements from visually implicit messages. We introduced three metrics—Creativity, Alignment, and Persuasiveness (CAP)—that achieved high agreement with human annotations. We showed existing T2I models struggle to generate images that are creative, aligned, and persuasive. We improved CAP scores through a simple yet effective approach using LLM-based expansion of abstract messages requiring interpretation of implicit content, and used these as prompts to T2I models.
Acknowledgment. This work was partly supported by NSF Grant No. 2006885 and partly by the University of Pittsburgh Center for Research Computing and Data, RRID:SCR_022735, through the resources provided. Specifically, this work used the H2P cluster, which is supported by NSF award number OAC-2117681. We gratefully acknowledge the support of our annotators.
@@
    Implementation details.
@@
    A detailed table for text-image alignment, including metrics from prior works, such as ImageReward, VQAScore and CLIPScore (Table 7).
  3. 3.
    Results from baseline metrics for evaluating creativity and persuasiveness (Table 8).
  4. 4.
    Two examples showing the benefit of combining the persuasiveness components for better alignment with human evaluations.
  5. 5.
@@
  6. 6.
    Example showing impact of different $\alpha$ values.
  7. 7.
@@
  9. 9.
    Human annotation interface.
7 Implementation Details
We fine-tuned LLAMA3-Instruct-8B (LLAMA3) and Qwen-2.5-7B-Instruct (QwenLM) LoRA [19] parameters, for generating accurate action-reason statements for each image ($AR_{gen}$), using the CPO trainer [52], with a batch size of 4, learning rate of 5e-5, and 3,000 training steps. We sampled 250 images, utilizing each of the three to five $AR_{m}$ messages from the PittAd dataset [21] as accepted statements in the training data. Each accepted statement was paired with all 4 hard negatives for the corresponding image from [30], to create negative statements. This process created a training set of 3,500 data points.
For LLAMA3-Instruct-8B [11], Qwen-2.5-7B-Instruct [2], QwenVL-2-7B-Instruct [3], and InternVL-2-26B [9], AuraFlow-2 [13] and FLUX.1-dev [15] we used the default temperature settings with 8-bit quantization. For SDXL [35] and FLUX.1-dev [15], we applied the default guidance scale, no negative prompt, and 28 inference steps. For AuraFlow-2 [13], we set a guidance scale of 5 and used 28 inference steps.
All experiments were conducted on A100 and L40s GPUs.
@@ 8 Alignment Results
Image COM Ads PSA Ads
$I_{input-text}$ T2I AIM IR VQA CS AIM IR VQA CS
@@
Table 7: Comparison of alignment of the images with action-reason message ($AR_{m}$) using our proposed AIM and baseline metrics. Best result per column bolded (including real images). IR = ImageReward. CS = CLIPScore. This result uses InternVL and LLAMA3.
Table 7 represents the text-image alignment scores for different images and evaluated by AIM and baseline metrics. Since $I_{real}$ are the ground-truth images in the dataset and $AR_{m}$ are the interpretation of real images, we expect $I_{real}$ to have higher alignment score than $I_{gen}$ (i.e. $I_{AR}$, $I_{LLM}$). However, the table highlights that ImageReward (IR) [53] and VQAScore (VQA) [28] assign lower scores to $I_{real}$ (ground-truth) compared to generated images. This contradicts our expectation and highlights the limitations of these metrics when evaluating the alignment of images and implicit prompts. CLIPScore (CS) [17] performs reasonably as it rates $I_{real}$ the highest, but it rates the alignment of $I_{LLM}$ and $I_{AR}$ equally. This is problematic because human annotators preferred $I_{LLM}$ in 92% of comparisons, demonstrating its superior alignment. AIM assigns the highest score to $I_{real}$, followed by $I_{LLM}$ which highlights the higher accuracy of our proposed evaluation method.
9 Creativity and Persuasiveness Baseline Metrics
Image COM Ads PSA Ads
$I_{input-text}$ T2I $C_{obj}$ $C_{LLM}$ $P_{c+A}$ $P_{LLM}$ $C_{obj}$ $C_{LLM}$ $P_{c+A}$ $P_{LLM}$
$I_{AR}$ SDXL 2.03 0.63 0.62 0.60 1.33 0.59 0.48 0.52
AuraFlow 2.12 0.73 0.64 0.60 1.36 0.72 0.42 0.57
$I_{LLM}$ AuraFlow 2.25 0.77 0.70 0.76 1.87 0.71 0.60 0.67
$I_{real}$ 2.28 0.67 0.98 0.66 2.04 0.65 0.60 0.61
Table 8: Comparison of persuasiveness and creativity proposed ($C_{obj}$, $P_{c+A}$) and baseline metrics. Best gen. image result bolded.
$I_{real}$ are images from PittAd dataset [21] designed to be creative and persuasive and are expected to be more creative than $I_{gen}$. As observed in Table 8 the $C_{LLM}$ score for $I_{real}$, which represents the ground-truth images, is lower than the scores for both $I_{LLM}$ and $I_{AR}$ generated by AuraFlow-2, by 10% and 6% respectively, across both Commercial and PSA advertisements. This raises questions about the reliability of LLMs in accurately scoring creativity, as real images receive lower scores than generated ones. In contrast, our proposed metrics assign the highest scores to $I_{real}$, demonstrating a more consistent alignment with human expectations.
10 Examples of Benefit of Combining Persuasiveness Components
Figure 6: Example of images chosen by human annotators for each component in $P_{comp}$ score compared to $P_{comp}$ score. In each row, a $\checkmark$ represents the chosen image for the corresponding component by human or the persuasiveness component, while a $\times$ indicates the rejected image. Green represents the agreement of score with human annotator while red represents disagreement.
@@ 11 Persuasiveness Metrics on Real Ads
Due to the inherent differences in quality between $I_{real}$ and $I_{gen}$, we avoided direct comparisons in human evaluations to reduce bias in selecting the better option. Instead, we compared two $I_{real}$ images from the same topic. Since the images were distinct, their corresponding $AR_{m}$ messages also differed, making alignment comparisons non-applicable. However, we evaluated the persuasiveness components for the two $I_{real}$ images and reported the agreement levels between the annotators, as well as between the annotators and the LLM, across all components. The results are presented in Table 9. We also analyze the agreement between the most frequently chosen images by annotators across all components and the average scores assigned to each image for these components. As observed in Table 9, while the agreement for individual components is lower than chance, combining all components results in an average agreement increase of 0.44. It is worth noting that the agreement between the $P_{comp+AIM}*$ metric and human annotators is higher than the agreement between the human annotators.
@@
Table 9: On real ads, agreement among annotators and the LLM responses for each persuasiveness component (column headers). “All” computes agreement across all components: for H, we get the image chosen most frequently across all the components, and for $P_{comp+AIM}$, we choose the image with higher average score.
11.1 Impact of Different $\alpha$ Values
Figure 7: Example of different values for alpha in the AIM score. In each row, the value under each image indicates the score generated by the AIM with the alpha listed for that row. A $\checkmark$ shows the chosen image, and a $\times$ indicates the rejected image. Green shows agreement with the human in choosing the more aligned image and red highlights the disagreement with the human.
Fig. 7 shows an example of AIM scores with different $\alpha$, in Eq. 1 in the main paper. Results are similar for values of $\alpha$ greater than one (i.e. greater impact of alignment over the reason in the ad message).
@@ 12 Failure Examples and Limitations
Figure 8: Example of failure of AIM in both generating accurate $AR_{gen}$ and choosing the more aligned image with $AR_{m}$. Red text indicates inaccurate action/reason considering the image. Green text shows accurate action/reason. Red circle shows inaccurate choice of image for which image is more aligned with the $AR_{m}$ according to the human annotation.
In this section, we first highlight examples of failures in the AIM method, specifically in generating accurate $AR_{gen}$ and selecting images that align with human annotators’ choices, as illustrated in Fig. 8. Next we discuss the limitation of our proposed metrics.
@@ 13 Prompts
In this section, we present the prompts used in different evaluation methods. We begin with the prompts for generating descriptions, as shown in Listings LABEL:prompt:train_description and LABEL:prompt:inference_description. Next, we provide the prompts for generating action-reason statements based on image descriptions in Listings LABEL:prompt:alignment_zeroshot and LABEL:prompt:alignment_ft. Then, we present the prompts used for evaluating the creativity and persuasiveness of images with LLMs, detailed in Listings LABEL:prompt:creativity and LABEL:prompt:persuasiveness. Finally, we outline the prompts for evaluating individual $P_{comp}$ components, including elaboration, synthesis, originality, imagination, audience targeting, benefit conversion, and appeal category, in Listings LABEL:prompt:elaboration, LABEL:prompt:synthesis, LABEL:prompt:originality, LABEL:prompt:imagination, LABEL:prompt:audience, LABEL:prompt:benefit, and LABEL:prompt:appeal.
@@
Listing 2: Description generation for images with MLLMs prompt templates for inference. If the answer to Q1 is “No”, then the AIM score for that image is 0.
@@
Listing 3: Prompt for Action-Reason Prediction with zero-shot LLAMA3 [11] and fine-tuned QwenLM [2]. {Blue} denotes elements added dynamically, based on the description of the image.
@@
Listing 4: Prompt for Action-Reason Prediction with fine-tuned LLM[11] in AIM evaluation. {Blue} denotes elements added dynamically, based on the description of the image.
@@
Listing 5: Prompt for evaluation of creativity with LLMs. {Blue} denotes elements added dynamically, based on the description of the image.
@@
Context: If the image convinces the audience to take an action, it is considered persuasive.
@@
Question: Based on the context score the persuasiveness of the described image in the range of (0, 5).
@@
Listing 7: Prompt for scoring elaboration with LLMs. {Blue} denotes elements added dynamically.
@@
You are given the description of an image and you are asked to score the image based on the following question.
@@
How visually detailed is the image? Do not consider the text in the image. Return a score in the range (0,5).
@@
Listing 8: Prompt for scoring synthesis with LLMs. {Blue} denotes elements added dynamically.
@@
You are given the description of an image and you are asked to score the image based on the following question.
@@
How well does the image connect the objects that are usually unrelated? Return the score in range (0,5).
@@
Listing 9: Prompt for scoring originality with LLM. {Blue} denotes elements added dynamically.
@@
You are given the description of an image and you are asked to score the image based on the following question.
@@
Listing 10: Prompt for scoring imagination with LLMs. {Blue} denotes elements added dynamically.
@@
You are given the description of an image and you are asked to score the image based on the following question.
@@
Assume you are a human evaluating the given image. How well does the image allow you to form images you have not directly experienced before more easily? Do not consider the text in the image. Return a score in the range (0,5)
@@
Listing 11: Prompt for scoring how well the image targets audience with LLMs. {Blue} denotes elements added dynamically.
@@
Listing 12: Prompt for scoring how well the image converts features to benefits for costumer with LLMs. {Blue} denotes elements added dynamically.
@@
You are given the description of an image and you are asked to score the image based on the following question.
@@
How well does the image connect the feature of the products to the benefits for customers? Return the score in range (0,5).
@@
Listing 13: Prompt for scoring how well the image targets correct appeal category with LLMs. {Blue} denotes elements added dynamically.
@@
You are given the description of an image and you are asked to score the image based on the following question.
@@
There are three types of rhetorical appeals or ways to convince the audience:
@@
Question: Assume you are a human evaluating the image. Based on the context score, how well the image appeals to {appealcategory} in the range of (0,5).